OpenStudy (anonymous):

Find these derivatives y=arcsec(x/3)

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

This should have everything you need: http://www.themathpage.com/acalc/inverse-trig.htm#arcsec

OpenStudy (anonymous):

ok but i got the answer \[\frac{ 1 }{ \left| x \right| \sqrt{x^2-9} }\]

OpenStudy (anonymous):

but the answer was \[\frac{ 3 }{ \left| x \right|\sqrt{x^2-9} }\]

OpenStudy (agent0smith):

You probably forgot to use the chain rule, with the derivative of the x/3 being 1/3.

OpenStudy (anonymous):

but arcsecu's derivative should be \[\frac{ u' }{ \left| u \right| \sqrt{u^2-1}}\]

OpenStudy (anonymous):

a=3 u=x u'=1

OpenStudy (anonymous):

so that's why i got \[\frac{ 1 }{ \left| x \right|\sqrt{x^2-9} }\] instead of \[\frac{ 3 }{ \left| x \right|\sqrt{x^2-9}}\]

OpenStudy (agent0smith):

If you're using arcsecu, then u=x/3, not u=x...

OpenStudy (anonymous):

so then my answer would be \[\frac{ 1 }{ \left| x \right| \sqrt{\frac{ x^2 }{ 9} - 9}}\]

OpenStudy (agent0smith):

You seem to be mixing things up... you just posted \[\frac{ u' }{ \left| u \right| \sqrt{u^2-1}}\]

OpenStudy (anonymous):

sorry top would be 1/3 instead of 1 rite?

OpenStudy (agent0smith):

u' = 1/3 u = x/3 u^2 = x^2/9

OpenStudy (agent0smith):

\[\huge \frac{ \frac{ 1 }{ 3 } }{ \left| \frac{ x }{ 3 } \right| \sqrt{\frac{ x^2 }{9 }-1}}\]

OpenStudy (agent0smith):

Now you can pull a 1/9 out of the square root...\[\huge \frac{ \frac{ 1 }{ 3 } }{ \left| \frac{ x }{ 3 } \right| \sqrt{\frac{ 1 }{ 9 }( x^2 -9)}} \]

OpenStudy (anonymous):

so how did i end up with 3 on top?

OpenStudy (anonymous):

the answer was \[\frac{ 3 }{ \left| x \right| \sqrt{x+1}}\]

OpenStudy (agent0smith):

Yes, take it from where I left off, simplify.

OpenStudy (anonymous):

x^2 + 9 i mean lol

OpenStudy (anonymous):

the answer was \[\frac{ 3 }{ \left| x \right|\sqrt{x^2-9} }\]

OpenStudy (anonymous):

idk how to simplify to get 3 on top

OpenStudy (agent0smith):

\[\huge \frac{ \frac{ 1 }{ 3 } }{ \left| \frac{ x }{ 3 } \right| \sqrt{\frac{ 1 }{ 9 }( x^2 -9)}}\] sqrt of 1/9 is 1/3, and dividing by a fraction is the same as multipling by it's reciprocal.

OpenStudy (anonymous):

oh wait nvm lol

OpenStudy (agent0smith):

haha :)

OpenStudy (anonymous):

can you help me with another one please? :)